OpenStudy (anonymous):

Question 2.2. Solve by the linear combination method (with or without multiplication). 2x + 3y = 11 –4x + 2y = 2 (Points : 1) (1, 3) (2, 3) (1, 2) (1, –2)

OpenStudy (amistre64):

what is the linear combination method?

OpenStudy (danjs):

multiply the first equation by -2 and then add them resulting in a cancellation of the x's

OpenStudy (danjs):

+2 i mean

OpenStudy (anonymous):

then wat?

OpenStudy (anonymous):

2x+3y=11 -4x+2y=2 2x+3y=11 -2x+y=1 -------- 4y=12 Y=3 Plug in -2x+3=1 -2x=-2 X=1 (1,3)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i have a few more

OpenStudy (danjs):

ok

OpenStudy (anonymous):



OpenStudy (anonymous):

?

OpenStudy (danjs):

U still have a question?

OpenStudy (anonymous):

BORDEM sorry

OpenStudy (anonymous):



OpenStudy (anonymous):

yea

OpenStudy (danjs):

Sorry this site is very slow, Multiply the first equation by 2, then add them to cancel the g's

OpenStudy (danjs):

-8g + 10h = 8 8g + 10h = -4 ---------------- 0 + 20h = 4 so h = 4/20 = 1/5

OpenStudy (danjs):

put h=1/5 into any of the equations to get the value of g.